dem0graphic profile of ecuator in 2018


27.08% population are 0-14 years old


18.35% population are 15-24 years old


39.59 population are 25-54 years old


7.53% population are 55-64 years old


7.45% population are 65 years old


number of random sample of 20 people .what is probability that 2 are 65 years ols, 5 are 55-64 years old, 6 are 25-54 years old ,the number of random sample that have composition like that one described above is________.Each of those random sample have a probability of occuring of _________the probability of event containing all random samples that fit the description above is______________


options


a) 97772875200


b)1.1649


c)0.0001138


d)349186


e)0.319


1
Expert's answer
2021-11-08T17:28:55-0500

To solve this question, we shall apply the multinomial distribution as described below. The multinomial distribution states that,

Let "c" denote the number of outcomes categories with their probabilities as, "\pi_1,\pi_2,...,\pi_c"

where "\displaystyle\sum_{i=1}^c(\pi_i)=1.0" then for "n" independent observations, the multinomial probabilities that "n_1" falls in category 1, "n_2" falls in category 2, ...., "n_c" falls in category c is given by

"p(n_1,n_2,..,n_c)=n!/(n_1!n_2!..n_c!)*\pi_1^{n_1}\pi_2^{n_2}*...*\pi_c^{n_c}"

For this question, "n=20,\pi_1=0.2708, \pi_2=0.1835,\pi_3=0.3959,\pi_4=0.0753,\pi_5=0.0745"

and "n_3=6,n_4=5,n_5=2"

Therefore, for this case, "n=6+5+2=13" and the probability is,

"p(n_3=6,n_4=5,n_5=2)=13!\/(6!5!2!)(0.3959^6)*(0.0753^5)*(0.0745^2)=1.407929e+13* 5.17369e-11=1.864391e-06"

Therefore, the probability that 2 are 65 years old, 5 are 55-64 years old, 6 are 25-54 years old is

The number of random sample that have composition like that one described above is 5, each of those random sample have a probability of occurring of "\pi_i", the probability of event containing all random samples that fit the description above is 1.0.